# The Complexity of Defining and Separating Fixpoint Formulae in Modal Logic

### v1

@@ 3. Foundations of Separability
We start with recalling the notion of separability and discuss some of its basic properties. {defi} Let $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$. A formula $\psi$ separates $\varphi$ and $\varphi^{\prime}$ if $\varphi\models\psi$ and $\psi\models\neg\varphi^{\prime}$. If $\psi$ comes from a fragment $\mathcal{L}$ of $\mu\mathsf{ML}$, then we call it an $\mathcal{L}$-separator of $\varphi$ and $\varphi^{\prime}$. It should be clear that a separator can only exist if $\varphi$ and $\varphi^{\prime}$ are inconsistent, and that either of $\varphi$ and $\neg\varphi^{\prime}$ is a trivial separator in this case. Thus, the notion is only interesting if we restrict $\mathcal{L}$. We illustrate the notion of a separator with a simple example. {exa} Consider the $\mu\mathsf{ML}$-formula $\theta_{\infty}=\nu x.\Diamond x$ expressing the existence of an infinite path originating in the root, and $\varphi^{\prime}=\Box\bot$ expressing that the root does not have any children. Clearly, $\theta_{\infty}$ and $\varphi^{\prime}$ are inconsistent, and for every $n\geq 1$ the modal formula $\Diamond^{n}\top$ separates $\theta_{\infty}$ and $\varphi^{\prime}$. Hence, there are infinitely many separators. Moreover, since $\Diamond^{m}\top\models\Diamond^{n}\top$ whenever $m\geq n$, there does not exist a logically strongest separator. ∎ The main problem we investigate in this paper is $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae, which asks whether given formulae $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$ admit an $\mathsf{ML}$-separator. Since any $\mathsf{ML}$-separator of $\varphi,\neg\varphi$ is equivalent to $\varphi$, a natural special case of $\mathsf{ML}$-separability is $\mathsf{ML}$-definability, in which we assume $\varphi^{\prime}=\neg\varphi$. The notion of separation is closely related to the notion of interpolation. Let $\mathcal{L}$ be a set of $\mu\mathsf{ML}$-formulae. A Craig interpolant of $\varphi,\varphi^{\prime}\in\mathcal{L}$ is an $\mathcal{L}$-separator $\psi$ of $\varphi,\neg\varphi^{\prime}$ which only uses the common symbols of $\varphi$ and $\varphi^{\prime}$, that is, $\text{sig}(\psi)\subseteq\text{sig}(\varphi)\cap\text{sig}(\varphi^{\prime})$. The Craig interpolant existence problem for $\mathcal{L}$ asks if a given pair of $\mathcal{L}$-formulae admit a Craig interpolant. We say that $\mathcal{L}$ enjoys the Craig interpolation property (CIP) if every $\varphi,\varphi^{\prime}\in\mathcal{L}$ with $\varphi\models\varphi^{\prime}$ admit a Craig interpolant. In logics enjoying the CIP, the Craig interpolant existence problem to checking if $\varphi\models\varphi^{\prime}$, but in logics which lack CIP the problem may me more involved.
@@
The starting point for the technical developments in the paper are model-theoretic characterizations for separability. Similar to what has been done in the context of interpolation, see for example [Rob60], they are given in terms of joint consistency, which we introduce next. Let $R$ be a binary relation on some class of models, such as $(\sigma,n)$-isomorphism $\cong_{\sigma}^{n}$ or $\mathsf{ML}_{\sigma}^{n}$-equivalence $\equiv_{\mathsf{ML}_{\sigma}^{n}}$. We call two formulae $\varphi,\varphi^{\prime}$ joint consistent up to $R$ (in short joint $R$-consistent) if there are models $\mathcal{M}\models\varphi$ and $\mathcal{M}^{\prime}\models\varphi^{\prime}$ with $R(\mathcal{M},\mathcal{M}^{\prime})$. For technical reasons we will sometimes also talk about joint consistency of automata $\mathcal{A},\mathcal{A}^{\prime}$ in place of formulae $\varphi,\varphi^{\prime}$. Joint $R$-consistency over a class $\mathbf{C}$ of models is defined by only looking at models from $\mathbf{C}$. Clearly, if $R^{\prime}\subseteq R$ and $\mathbf{C}^{\prime}\subseteq\mathbf{C}$ then joint $R^{\prime}$-consistency over $\mathbf{C}^{\prime}$ implies joint $R$-consistency over $\mathbf{C}$. We use the following standard equivalence:
@@ 3.1. Modal Definability: A Recap
In his seminal paper [Ott99] Otto showed that $\mathsf{ML}$-definability of $\mu\mathsf{ML}$-formulae is ExpTime-complete over all models and over $\mathbb{T}^{d}$ for every $d\geq 2$. {thmC}[Ott99, Main Theorem and Proposition 5] Over the class of all models, as well as over $\mathbb{T}^{d}$ for every $d\geq 2$, $\mathsf{ML}$-definability of $\mu\mathsf{ML}$-formulae is ExpTime-complete. We start by recalling and rephrasing Otto’s construction and fixing a small mistake in the original proof. The lower bound follows by an immediate reduction from satisfiability of $\mu\mathsf{ML}$-formulae. We look at the upper bound. The first step is the following lemma, which is the heart of [Ott99, Lemma 2].
@@
The lemma is true, but its proof in [Ott99] is mistaken. The problem there is that the construction duplicates subtrees and hence may turn $d$-ary models into ones with outdegree greater than $d$. We present an easy alternative proof.
@@
Using automata-based techniques we can decide if Item 2 in Lemma 3 holds for all $n$.
@@
Let $\mathcal{A},\mathcal{A}^{\prime}$ be parity automata and $d\in\mathbb{N}$. Then, $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ for all $n\in\mathbb{N}$ iff $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\cong_{\sigma}^{m}$-consistent over $\mathbb{T}^{d}$ for $m=|\mathcal{A}|\times|\mathcal{A}^{\prime}|+1$. The latter condition can be checked in time polynomial in $|\mathcal{A}|+|\mathcal{A}^{\prime}|$.
@@
(Sketch) Over $\mathbb{T}^{d}$, $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent iff they are joint $\cong_{\sigma_{0}}^{n}$-consistent for $\sigma_{0}=\sigma\cap(\text{sig}(\mathcal{A})\cup\text{sig}(\mathcal{A}^{\prime}))$. Thus, without loss of generality we assume that $\sigma\subseteq\text{sig}(\mathcal{A})=\text{sig}(\mathcal{A}^{\prime})$. Due to well-known relativization techniques we do not loose generality by only running $\mathcal{A},\mathcal{A}^{\prime}$ on full $d$-ary trees with no leaves. Let $L$ be the language of finite full $d$-ary trees over $\sigma$ such that $\mathcal{M}\in L$ iff $\mathcal{M}$ is a prefix of a $\sigma$-reduct of a model of $\mathcal{A}$. Let $L^{\prime}$ be an analogous language for $\mathcal{A}^{\prime}$. The tallness of a finite tree is the minimal distance from the root to a leaf. Observe that $\mathcal{A},\mathcal{A}^{\prime}$ are $\cong_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ iff $L\cap L^{\prime}$ contains a tree of tallness $n$. Thus, it suffices to check if $L\cap L^{\prime}$ contains trees of arbitrarily high tallness. To that end construct an automaton $\mathcal{B}$ recognizing $L\cap L^{\prime}$ of size $|\mathcal{A}|\times|\mathcal{A}^{\prime}|$. An easy pumping argument shows that the language $L\cap L^{\prime}$ of $\mathcal{B}$ contains trees of arbitrarily high tallness iff it contains a tree of tallness $m=|\mathcal{B}|+1$. To test the latter condition it is enough to inductively compute a sequence $S_{1}\supseteq S_{2}\supseteq...\supseteq S_{|\mathcal{B}|+1}$ of subsets of states of $\mathcal{B}$, where $S_{i}$ is the set of all states $q$ such that $\mathcal{B}[q_{I}\mapsfrom q]$ recognizes a tree of tallness at least $i$.
@@
We are ready to solve $\mathsf{ML}$-definability over $\mathbb{T}^{d}$ in exponential time. Assume $\mu\mathsf{ML}$-formula $\varphi$ over signature $\sigma$. For every $n$, we know by Equivalence ($\mathsf{Base}$) that $\varphi$ is equivalent over $\mathbb{T}^{d}$ to some $\psi\in\mathsf{ML}_{\sigma}^{n}$ iff $\varphi,\neg\varphi$ are not joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$. By Lemma 3 this is equivalent to the lack of joint $\cong_{\sigma}^{n}$-consistency of $\varphi,\neg\varphi$ over $\mathbb{T}^{d}$. By Theorem 1 we can compute exponentially-sized automata $\mathcal{A}$, $\mathcal{A}^{\prime}$ equivalent to $\varphi$ and $\neg\varphi$ over $\mathbb{T}^{d}$. It follows that $\varphi$ is not $\mathsf{ML}$-definable over $\mathbb{T}^{d}$ iff $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ for every $n$. The last condition is decided using Proposition 4. The runtime of our algorithm is polynomial in $|\mathcal{A}|+|\mathcal{A}^{\prime}|$, and thus exponential in $|\varphi|$. This proves the part of Theorem 3.1 about $\mathbb{T}^{d}$. The remaining part concerning unrestricted models is a special case of Theorem 6, which we will prove in the following subsection.
@@ 3.2. Modal Separation: the Unrestricted Case
Let $\mathbf{C}$ be the class of all models or $\mathbb{T}^{2}$. Assume an NPTA $\mathcal{A}$ over $\mathbf{C}$, and in the case $\mathbf{C}=\mathbb{T}^{2}$ assume that $\mathcal{A}$ is duplication safe. For every $n\in\mathbb{N}$, an $\mathsf{ML}^{n}$-uniform consequence of $\mathcal{A}$ over $\mathbf{C}$ can be constructed in time $|\mathcal{A}|^{O(n\cdot|\mathcal{A}|)}$ if $\mathbf{C}$ is the class of all models and in time $2^{O(n\cdot|\mathcal{A}|)}$ if $\mathbf{C}=\mathbb{T}^{2}$.
@@
Let $\mathcal{A}=(Q,\Sigma,q_{I},\delta,\mathsf{rank})$ be an NPTA over $\mathbf{C}$. In the case $\mathbf{C}=\mathbb{T}^{2}$ assume that $\mathcal{A}$ is duplication safe. We construct $\psi_{m,q}$ for every $q\in Q$ and $m\leq n$ by induction on $m$. For the base case we put:
@@
for every structure $\mathcal{M}\in\mathbf{C}$. Hence, $\psi_{n,q_{I}}$ is an $\mathsf{ML}^{n}$-uniform consequence of $\mathcal{A}$ over $\mathbf{C}$. It is routine to check that the formula $\psi_{n,q_{I}}$ has the right size.
@@
It is not difficult to show that, in the presence of at least two accessibility relations $\Diamond_{1},\Diamond_{2}$, the construction is optimal: one can express in $\mu\mathsf{ML}$ that the model embeds a full binary tree of depth $2^{n}$ and in which each inner node has both a $\Diamond_{1}$- and a $\Diamond_{2}$-successor. Using standard techniques, one can show that any modal formula expressing this property is of doubly exponential size [FvdHIK13]. Whether having two accessibility relations is necessary for this lower bound is an interesting question which we leave open.

### v2

@@ 3. Foundations of Separability
We start with recalling the notion of separability and discuss some of its basic properties. {defi} Let $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$. A formula $\psi$ separates $\varphi$ and $\varphi^{\prime}$ if $\varphi\models\psi$ and $\psi\models\neg\varphi^{\prime}$. If $\psi$ comes from a fragment $\mathcal{L}$ of $\mu\mathsf{ML}$, then we call it an $\mathcal{L}$-separator of $\varphi$ and $\varphi^{\prime}$. It should be clear that a separator can only exist if $\varphi$ and $\varphi^{\prime}$ are inconsistent, and that either of $\varphi$ and $\neg\varphi^{\prime}$ is a trivial separator in this case. Thus, the notion is only interesting if we restrict $\mathcal{L}$. We illustrate the notion of a separator with a simple example. {exa} Consider the $\mu\mathsf{ML}$-formula $\theta_{\infty}=\nu x.\Diamond x$ expressing the existence of an infinite path originating in the root, and $\varphi^{\prime}=\Box\bot$ expressing that the root does not have any children. Clearly, $\theta_{\infty}$ and $\varphi^{\prime}$ are inconsistent, and for every $n\geq 1$ the modal formula $\Diamond^{n}\top$ separates $\theta_{\infty}$ and $\varphi^{\prime}$. Hence, there are infinitely many separators. Moreover, since $\Diamond^{m}\top\models\Diamond^{n}\top$ whenever $m\geq n$, there does not exist a logically strongest separator. ∎ The main problem we investigate in this paper is $\mathsf{ML}$-separability of $\mu\mathsf{ML}$-formulae, which asks whether given formulae $\varphi,\varphi^{\prime}\in\mu\mathsf{ML}$ admit an $\mathsf{ML}$-separator. Since any $\mathsf{ML}$-separator of $\varphi,\neg\varphi$ is equivalent to $\varphi$, a natural special case of $\mathsf{ML}$-separability is $\mathsf{ML}$-definability, in which we assume $\varphi^{\prime}=\neg\varphi$. The notion of separation is closely related to the notion of interpolation. Let $\mathcal{L}$ be a set of $\mu\mathsf{ML}$-formulae. A Craig interpolant of $\varphi,\varphi^{\prime}\in\mathcal{L}$ is an $\mathcal{L}$-separator $\psi$ of $\varphi,\neg\varphi^{\prime}$ which only uses the common symbols of $\varphi$ and $\varphi^{\prime}$, that is, $\text{sig}(\psi)\subseteq\text{sig}(\varphi)\cap\text{sig}(\varphi^{\prime})$. The Craig interpolant existence problem for $\mathcal{L}$ asks if a given pair of $\mathcal{L}$-formulae admit a Craig interpolant. We say that $\mathcal{L}$ enjoys the Craig interpolation property (CIP) if every $\varphi,\varphi^{\prime}\in\mathcal{L}$ with $\varphi\models\varphi^{\prime}$ admit a Craig interpolant. In logics enjoying the CIP, the Craig interpolant existence problem reduces to checking if $\varphi\models\varphi^{\prime}$, but in logics which lack CIP the problem may me more involved.
@@
The starting point for the technical developments in the paper are model-theoretic characterizations for separability. Similar to what has been done in the context of interpolation, see for example [34], they are given in terms of joint consistency, which we introduce next. Let $R$ be a binary relation on some class of models, such as $(\sigma,n)$-isomorphism $\cong_{\sigma}^{n}$ or $\mathsf{ML}_{\sigma}^{n}$-equivalence $\equiv_{\mathsf{ML}_{\sigma}^{n}}$. We call two formulae $\varphi,\varphi^{\prime}$ joint consistent up to $R$ (in short joint $R$-consistent) if there are models $\mathcal{M}\models\varphi$ and $\mathcal{M}^{\prime}\models\varphi^{\prime}$ with $R(\mathcal{M},\mathcal{M}^{\prime})$. For technical reasons we will sometimes also talk about joint consistency of automata $\mathcal{A},\mathcal{A}^{\prime}$ in place of formulae $\varphi,\varphi^{\prime}$. Joint $R$-consistency over a class $\mathbf{C}$ of models is defined by only looking at models from $\mathbf{C}$. Clearly, if $R^{\prime}\subseteq R$ and $\mathbf{C}^{\prime}\subseteq\mathbf{C}$ then joint $R^{\prime}$-consistency over $\mathbf{C}^{\prime}$ implies joint $R$-consistency over $\mathbf{C}$. We use the following standard equivalence:
@@ 3.1. Modal Definability: A Recap
In his seminal paper [30] Otto showed that $\mathsf{ML}$-definability of $\mu\mathsf{ML}$-formulae is ExpTime-complete over all models and over $\mathbb{T}^{d}$ for every $d\geq 2$. {thmC}[30, Main Theorem and Proposition 5] Over the class of all models, as well as over $\mathbb{T}^{d}$ for every $d\geq 2$, $\mathsf{ML}$-definability of $\mu\mathsf{ML}$-formulae is ExpTime-complete. We start by recalling and rephrasing Otto’s construction and fixing a small mistake in the original proof. The lower bound follows by an immediate reduction from satisfiability of $\mu\mathsf{ML}$-formulae. We look at the upper bound. The first step is the following lemma, which is the heart of [30, Lemma 2].
@@
The lemma is true, but its proof in [30] is mistaken. The problem there is that the construction duplicates subtrees and hence may turn $d$-ary models into ones with outdegree greater than $d$. We present an easy alternative proof.
@@
Using automata-based techniques we can decide whether Item 2 in Lemma 3 holds for all $n$.
@@
Let $d\in\mathbb{N}$ and $\mathcal{A},\mathcal{A}^{\prime}$ be parity automata over $d$-ary trees with respective alphabets $\Sigma,\Sigma^{\prime}$. Then, $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ for all $n\in\mathbb{N}$ iff $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\cong_{\sigma}^{m}$-consistent over $\mathbb{T}^{d}$ for $m=|\mathcal{A}|\times|\mathcal{A}^{\prime}|+1$. If $\mathcal{A},\mathcal{A}^{\prime}$ are in trash normal form then the latter condition can be checked in time polynomial in $(|\Sigma|+|\Sigma^{\prime}|)\times(|\mathcal{A}|+|\mathcal{A}^{\prime}|)^{d}$.
@@
(Sketch) Without loss of generality both $\mathcal{A}$ and $\mathcal{A}^{\prime}$ work over the same alphabet, $\Sigma=\Sigma^{\prime}=\mathcal{P}(\sigma_{\textsf{aut}})$ for some signature $\sigma_{\textsf{aut}}$, and $\sigma\subseteq\sigma_{\mathsf{aut}}$. Due to well-known relativization techniques we also do not lose generality by only running $\mathcal{A},\mathcal{A}^{\prime}$ on full $d$-ary trees with no leaves. Let $L$ be the language of finite full $d$-ary trees over $\sigma$ such that $\mathcal{M}\in L$ iff $\mathcal{M}$ is a prefix of a $\sigma$-reduct of a model of $\mathcal{A}$. Let $L^{\prime}$ be the analogous language for $\mathcal{A}^{\prime}$. The tallness of a finite tree is the minimal distance from the root to a leaf. Observe that $\mathcal{A},\mathcal{A}^{\prime}$ are $\cong_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ iff $L\cap L^{\prime}$ contains a tree of tallness $n$. Thus, it suffices to check if $L\cap L^{\prime}$ contains trees of arbitrarily high tallness.
To that end construct an automaton $\mathcal{B}_{0}$ recognizing $L$: $\mathcal{B}_{0}$ has the same set of states as $\mathcal{A}$, and is computed in time polynomial in $|\mathcal{A}|$ and $|\Sigma|$ thanks to the trash normal form of $\mathcal{A}$. Let $\mathcal{B}_{0}^{\prime}$ be an analogous automaton for $L^{\prime}$. Construct from $\mathcal{B}_{0},\mathcal{B}_{0}^{\prime}$ an automaton $\mathcal{B}$ recognizing $L\cap L^{\prime}$, of size $|\mathcal{B}|=|\mathcal{B}_{0}|\times|\mathcal{B}_{0}^{\prime}|=|\mathcal{A}|\times|\mathcal{A}^{\prime}|$. An easy pumping argument shows that the language $L\cap L^{\prime}$ of $\mathcal{B}$ contains trees of arbitrarily high tallness iff it contains a tree of tallness $m=|\mathcal{B}|+1$. To test the latter condition it is enough to inductively compute a sequence $S_{1}\supseteq S_{2}\supseteq...\supseteq S_{|\mathcal{B}|+1}$ of subsets of states of $\mathcal{B}$, where $S_{i}$ is the set of all states $q$ such that $\mathcal{B}[q_{I}\mapsfrom q]$ recognizes a tree of tallness at least $i$.
@@
We are ready to solve $\mathsf{ML}$-definability over $\mathbb{T}^{d}$ in exponential time. Assume $\mu\mathsf{ML}$-formula $\varphi$ over signature $\sigma$. For every $n$, we know by Equivalence ($\mathsf{Base}$) that $\varphi$ is equivalent over $\mathbb{T}^{d}$ to some $\psi\in\mathsf{ML}_{\sigma}^{n}$ iff $\varphi,\neg\varphi$ are not joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$. By Lemma 3 this is equivalent to the lack of joint $\cong_{\sigma}^{n}$-consistency of $\varphi,\neg\varphi$ over $\mathbb{T}^{d}$. By Theorem 1 we can compute exponentially-sized automata $\mathcal{A}$, $\mathcal{A}^{\prime}$ in trash normal form, equivalent to $\varphi$ and $\neg\varphi$ over $\mathbb{T}^{d}$. It follows that $\varphi$ is not $\mathsf{ML}$-definable over $\mathbb{T}^{d}$ iff $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ for every $n$. The last condition is decided using Proposition 4. The runtime of our algorithm is polynomial in $(|\Sigma|+|\Sigma^{\prime}|)\times(|\mathcal{A}|+|\mathcal{A}^{\prime}|)^{d}$, and thus exponential in $|\varphi|$. This proves the part of Theorem 3.1 about $\mathbb{T}^{d}$. The remaining part concerning unrestricted models is a special case of Theorem 6, which we will prove in the following subsection.
@@ 3.2. Modal Separation: the Unrestricted Case
Let $\mathbf{C}$ be the class of all models or $\mathbb{T}^{2}$, and let $\mathcal{A}$ be an NPTA over $\mathbf{C}$ with alphabet $\Sigma$. Assume that $\mathcal{A}$ is in the trash normal form, and in the case $\mathbf{C}=\mathbb{T}^{2}$ also that it is duplication safe. For every $n\in\mathbb{N}$, an $\mathsf{ML}^{n}$-uniform consequence of $\mathcal{A}$ over $\mathbf{C}$ can be constructed in time $(|\Sigma|\times 2^{\mathcal{A}})^{O(n)}$} if $\mathbf{C}$ is the class of all models and in time $(|\Sigma|\times|\mathcal{A}|)^{O(n)}$ if $\mathbf{C}=\mathbb{T}^{2}$.
@@
Let $\mathcal{A}=(Q,\Sigma,q_{I},\delta,\mathsf{rank})$ be an NPTA over $\mathbf{C}$ in trash normal form. In the case $\mathbf{C}=\mathbb{T}^{2}$ assume that $\mathcal{A}$ is also duplication safe. We construct $\psi_{m,q}$ for every $q\in Q$ and $m\leq n$ by induction on $m$. For the base case we put:
@@
for every structure $\mathcal{M}\in\mathbf{C}$. Hence, $\psi_{n,q_{I}}$ is an $\mathsf{ML}^{n}$-uniform consequence of $\mathcal{A}$ over $\mathbf{C}$. Thanks to the trash normal form of $\mathcal{A}$, the base case formula is computed in polynomial time. The inductive case consists of a simple syntactic manipulation. Thus, the claimed complexity follows from the fact that $\psi_{q,n}$ has the right size for all $q$ and $n$, which can be checked by a routine induction on $n$.
@@
It is not difficult to show that, in the presence of at least two accessibility relations $\Diamond_{1},\Diamond_{2}$, the construction is optimal: one can express in $\mu\mathsf{ML}$ that the model embeds a full binary tree of depth $2^{n}$ and in which each inner node has both a $\Diamond_{1}$- and a $\Diamond_{2}$-successor. Using standard techniques, one can show that any modal formula expressing this property is of doubly exponential size [16]. Whether having two accessibility relations is necessary for this lower bound is an interesting question which we leave open.